From the known mass and radius of the moon, compute the value of the acceleration due to gravity, gM at the surface of the moon.


(mass of the moon M = 7.35 X10²² kg) ( G = 6.67 X10⁻¹¹ Nm²/kg² )


(radius of the moon R = 1.7374 X10⁶ )



1
Expert's answer
2022-03-07T12:21:29-0500

Answer

the value of the acceleration due to gravity

"g=\frac{GM}{R^2}\\=\frac{6.67*10^{-11}*7.35*10^{22}}{(64*10^5)^2}\\=0.126m/s^2"